OpenStudy (anonymous):

I could really use some help with this integral using trig identities. I'm terrible with trig. integral of x^2/((9-x^2)^(1/2))

OpenStudy (unklerhaukus):

\[\int\frac{x^2}{\sqrt{9-x^2}}\mathrm dx\]

OpenStudy (unklerhaukus):

the integral corresponds to the first case in the table , so substitute \(x=\frac ba\sin\theta\) . . .

OpenStudy (anonymous):

ok...I'm not really sure where to go from there with it though..